# VQE for Ising Model & A Comparative Analysis of Classical and Quantum Optimization Methods

### v1

@@ III.1 Transverse Ising model (TIM)
The Transverse Ising model was introduced first in 1963 by de Gennes when he built a model for describing the low-frequency collective modes of protons in the ferroelectric phase of ferroelectric crystals $\text{KH}_{2}\text{PO}_{4}$ [15]. From then until now, the Transverse Ising Model has been used to represent enormous simple-to-complex systems such as brain science and information science and technology.
@@
One of the key differences between the classical and quantum versions of the Ising model is that the expectation value of two terms in the Eq. (4) must be measured separately due to the non-commutative feature of these elements. However, the commutation property allows us to perform only two observations to extract the Hamiltonian’s expectation. The first term $\sum_{n=1}^{N-1}\sigma^{z}_{n-1}\sigma^{z}_{n}$ can be cumulatively achieved by measuring directly on the computational basis as follows:
@@ III.2 Ansatz Construction
III.2.1 Symmetry of the Transverse Ising model
@@
  • Real representation. Using the eigenstates of the $\sigma_{Z}$ (Pauli-z) operator as the elementary binary computational basis, in terms of which, $\sigma^{z}$, $\sigma^{x}$ are real matrices, due to that, we can represent the TIM Hamiltonian in real form, the eigenstates of the TIM Hamiltonian can thus be chosen to be real for conventional purposes. Namely, considering $|\Psi\rangle=\sum_{n}C_{n}|n\rangle$ is an eigenstate of $H_{\text{TIM}}$ expanding in the computational basis $|n\rangle$. The real form of $H_{\text{TIM}}$ means the Hermitian real element matrix, that induces coefficients satisfied $C^{*}_{n}C_{m}=C_{n}C^{*}_{m}$ $\forall m,n\in[0,2^{N}-1]$, then, in generality
@@
    The interaction of two local spins formulated by this term causes a sort of entanglement structure between every two local spins of the ground state energy. Nevertheless, in the case of the order phase, when the external magnetic field is dominant, this interaction can be broken down, each spin will be free interacting and aligns in the same direction as the magnetic field.
  • Total spin-flip symmetry . The most featured symmetry of the Ising model, can be referred to as the classical counterpart, Time-reversal symmetry (generally, $\mathcal{Z}_{2}$ symmetry). Under the total spin-flip transformation in the z direction $\left(\sigma^{x}\right)^{\otimes N}$
@@
III.2.2 Ansatz selection
For the chosen ansatz, the conventional real coefficients of eigenstates tell us that it is enough to span in real quantum parameter space for finding the ground state energy, and the linear entanglement mapping comes from the information of the local interaction term in the Hamiltonian. Moreover, to accommodate the device constraints, in a speculative way, the available common-use RealAmplitudes ansatz is our good candidate for the implementation, which uses only the quantum gate $R_{y}\left(\bf\theta\right)$ for the parametrized quantum circuit and the entanglement scheme as in Fig. 2.
Figure 2: A RealAmplitudes circuit with 2 layers on 3 qubits and the linear entanglement scheme.
The additional reason for choosing such a Hardware-efficient ansatz architecture [16] comes from the efficient implementation of quantum optimization to attain our full quantum algorithm purpose, which will be addressed in the next subsection IV.2.
To efficiently travel around full quantum Hilbert space, we need to tune $22^{N+1}-$ 2 degrees of freedom, $+1$ inside of the exponent indicates the complex number condition, and $-2$ turns up from normalization and global phase conditions. After bringing TIM’s properties we investigated from section III.2.1 to dissect, Real representation Eq. (7) helps us eliminate $+1$ within the exponent and Total spin-flip symmetry Eq. (11) reduces the number of parameters we have to search the ground state by half. Finally,
@@ III.2.2 Ansatz selection
The number of layers $L$, which changes directly ansatz number of parameters $p$ according to $p=N(L+1)$. Following Eq. (12), we roughly evaluate a necessary number of layer
@@
notes that this is not a rigorous condition because ansatz’s entanglement mapping scheme used also contributes, which is able to change the bound as well, it is still a good estimation for surface analysis nonetheless. Yet, we can realize that, not only TIM, but Eq. (12) also holds for all symmetrically real Hamiltonians being symmetric under $\left(\sigma^{x}\right)^{\otimes N}$ the operator. Therefore, there are some hidden TIM properties that we are unable to analytically incorporate to give an exact boundary value of $L$. Through conducting the experimental survey though, we are able to choose a reliable value, as we will see in chapter V.2, it works well even for choosing $L=2$ for all qubit scaling.
IV Optimization
The next subroutine in the VQE work map, which is a purely technical perspective we investigate to study the performance of different optimization methods, provides us a larger view of the way to reaching the faster and more precise convergence of our target value. In TIM, we try to minimize the cost function defined as the expectation value of Hamiltonian $\langle\hat{H}_{\text{TIM}}\rangle$ using common methods of optimization categorized into two class operations.
IV.1 Classical Operation
@@
Constrained Optimization BY Linear Approximation (COBYLA). One of the most powerful derivative-free methods, which has been favored by many users in lately decades, COBYLA makes use of linear interpolation of the objective function at each iteration by a unique linear polynomial function at the vertices for finding an optimal vector parameter within the trust region, then feeding optimal point evaluated to the objective function to get the value improving the next iteration of approximation [17, 18, 19, 20]. In VQE, COBYLA’s linear interpolation at each step is a classical subroutine running on the CPU using only one objective value from the QPU evaluation running. And so because of ignored derivative information, it would beneficially avoid several problems in analysis optimizations, especially Barren plateaus landscape, in return less accuracy for more parameters ($p>9$) [20]. For a bird’s-eye view, new updates to COBYLA versions like UOBYQA, NEWUOA, and BOBYQA count the objective function’s curvature information to increase convergence [21, 22, 23].
Finite Difference (FD). Let’s dawn on one of the primary first-order derivative-based approaches, where the parameter ${\bm{\theta}}\in\mathbb{R}^{p}$ is updated at the iteration $k$-th by
@@
where $\vec{i}\in\mathbb{R}^{p}$ is $i$-th the unit vector and $\epsilon$ is an infinitesimal change, with the error proportional to $\mathcal{O}(\epsilon^{2})$. The smaller $\epsilon$ we use, the more exact result we get. However, due to the restricted accuracy of a classical computer, we are unable to achieve a value of $\epsilon$ that is too small, which even becomes greater when implemented on a quantum device, where at least the sample error enters the picture.
Simultaneous Perturbation Stochastic Approximation (SPSA). To surmount some of obstacles that emerge from gradient descent operation on the near-term devices, the idea of a perturbing stochastic approximation method is a good choice. SPSA generates an unbiased estimator $\tilde{f}({\bm{\theta}}_{k})$ of the gradient by simultaneously randomly perturbating the gradient direction of parameters [24], we replace the ordinary gradient vector $\triangledown f({\bm{\theta}}_{k})$ by
@@
where $\vec{\Delta}_{k}$ is a random perturbed vector sampled from the zero-mean distribution, usually, the Bernoulli distribution is used. We can see that all parameters are simultaneously shifted by a random amount (±$\pm s_{k}$), our computation therefore only requires two objective function evaluations per iteration. Whereas standard gradient computation time is scaled along to the number of parameters, the SPSA optimizer is, however, independent, which saves a lot of time as we work on a higher parameter regime. Besides that, noise from quantum circuit executions computing the objective value be regarded as a stochastic perturbation part absorbed into the SPSA procedure. These advantages promote SPSA and its other versions to be efficient techniques in the NISQ era.
IV.2 Quantum Operation
As opposed to the classical ones, the quantum operations require the quantum information extracted from the structure of the trial wave function. This sort of behavior inevitably entails the optimization process in whole problem analysis, especially related closely to Ansatz Construction.
Parameter-shift rules (PSR). Inspired by the classical shift rules in the exact derivative computation of some special function, we are successfully able to figure out the analytical value of the objective function’s derivatives using quantum devices. For any kind of quantum gate has the form
@@ IV.2 Quantum Operation
generated by the Hermitian operator $\hat{G}$ Supposed that we have an expectation value of $\hat{H}$, which is our objective function $f(\theta)=\langle\psi(\theta)|\hat{H}|\psi(\theta)\rangle$ parametrized by $\theta$. The ansatz wave function $|\psi(\theta)\rangle=\hat{U}(\theta)|\psi\rangle_{I}$ is made up of $\hat{U}(\theta)=\hat{A}\hat{\mathcal{G}}(\theta)\hat{B}$, with $\hat{A}$, $\hat{B}$ are arbitrarily other operators. As a result, the partial derivative of the $f(\theta)$ with respect to $\theta$ is thus obtained via the parameter-shift rules [25]
@@
The partial derivative $\partial_{\theta}=\frac{\partial}{\partial\theta}$ implies we can generalize to a set of multiple parameters $\{\theta_{i}\}$. The evaluation of $f(\theta\pm\frac{\pi}{4s})$ can be easy to run on quantum computers, and then the exact amplitude of the vector gradient is derived. Within the aim of this paper, the given parametric ansatz is made out of Pauli rotations, and $s=\frac{1}{2}$ is selected accordingly. The state-of-the-art formalism for more generic PQCs is also able to be derived [26, 27].
Quantum Natural Gradient Descent (QNG). How to improve the convergence? The global fixed learning rate $\eta_{k}$ is apparently not a good choice, one does not be good sensitive to the model information with respect to parameter changes. Initially, many attempts try to tune the learning rate $\eta_{k}$ such as learning rate schedulers that vary $\eta_{k}$ after an iteration, latterly, adaptive methods that count previous iteration values, or capturing the curvature of the objective function using the diagonal approximation of the Hessian (HES), where each $\eta_{k}$ is the inverse Hessian’s diagonal element instead of being equally fixed. Those methods educe that we can embed further information rather than only just gradient vector, to make the optimal step size for our variational quantum algorithms. Likewise, the classical counterpart using the Fisher Information Matrix (FIM), Quantum Natural Gradient Descent invokes the quantum geometry of the wave function, which can make us get the optimal spot faster. Namely, we transform to the Riemann parameter space using the metric $g\in\mathbb{R}^{p\times p}$ of projected Hilbert space $\mathcal{PH}$ instead of the current Euclidean space $g=\mathbb{1}_{p\times p}$. The update Eq. (14) turns to
@@
We expand $|\psi_{\theta+d\theta}\rangle=|\psi_{\theta}\rangle+|\partial_{i}\psi_{\theta}\rangle d\theta_{i}+\frac{1}{2}|\partial_{i}\partial_{j}\psi_{\theta}\rangle d\theta_{i}d\theta_{j}$ upto second order in $d\theta$, where $\partial_{i}=\frac{\partial}{\partial\theta_{i}}$. The Quantum Geometry tensor (QGT) $G_{ij}\in\mathbb{R}^{p\times p}$ has two parts: the anti-symmetric imaginary part $\sigma_{ij}=-\sigma_{ji}$ related to the gauge field of $U(1)$ eventually vanishes, the only contribution comes from the real symmetric part $g_{ij}\equiv g_{ij}({\bm{\theta}})$ reflects the quantum distance in $\mathcal{PH}$ space, where its formula is
@@
  • QNG - Block Diagonal Approximation (QN-BDA). Moreover, more than Diagonal Approximation, where we just count $p$ elements in the diagonal line, the QN-BDA employing the Parametric Family circuit is conveniently deployed entirely in the computation of the approximated metric tensor on the quantum computer. The parametric unitary operator $\hat{U}({\bm{\theta}})$ acts on the initial state $|\psi\rangle_{I}$ entailing $L$ layers
@@
    • $S_{l}$ are the static parts, which usually denote to the entanglement layer.
@@
    • The single qubit gate $R_{i}(\theta^{l}_{i})=\exp{\left[i\theta_{i}^{l}K_{i}\right]}$ ) is constructed from Hermitian generator $K_{i}$ with the parameter $\theta_{i}^{l}\in\theta^{l}=\left\{\theta^{l}_{1},\ldots,\theta^{l}_{N}\right\}$.
Such a type of parametrized circuit whose nice properties we can prospect to yield a block diagonal form of QGT running completely on the quantum processor, each block corresponding to each layered vector parameter $\theta^{l}\in{\bm{\theta}}=\theta^{1}\oplus\ldots\oplus\theta^{L}$. We denote
@@
$|\partial_{i}\psi_{\theta}\rangle$ $=\partial_{i}\hat{U}({\bm{\theta}})|\psi\rangle_{I}$ (25)
$=\hat{U}^{L}_{l+1}S_{l}\partial_{i}P_{l}(\theta^{l})\hat{U}^{l-1}_{1}|\psi\rangle_{I}$ (26)
$=\hat{U}^{L}_{l}\left(iK_{i}\right)\hat{U}^{l-1}_{1}|\psi\rangle_{I}$ (27)
$=\hat{U}^{L}_{l}\left(iK_{i}\right)|\psi^{(l-1)}_{\theta}\rangle.$ (28)
Note that $\left[K_{i},K_{j}\right]=0$ in each layer, the unitarity of $\hat{U}^{m}_{n}$ and overlap of partial derivative states tell us the block matrix element
@@
Quantum Natural-Simultaneous Perturbation Stochastic Approximation (QN-SPSA). Inheriting the SPSA idea to calculate the gradient, it is capable to put up generating the Hessian matrix with fewer evaluations, the so-call 2-SPSA. Normally, the HES or FIM matrix consumes $\mathcal{O}(p^{2})$ quantum expectation, QN-BDA enables us to turn to relying on the number of layers $\mathcal{O}(L)$, saving us lots of computational sources. Even further, QN-SPSA just only executes four quantum runs to obtain the full second-order derivative matrix. Let’s consider the Hessian of the Fubini-Study metric,
@@
which is just the Hessian form of Eq. (22) so that we can deploy the 2-SPSA method to generate the QN-SPSA matrix. We can see the equivalence
@@
$=-\text{Re}\left\{\langle\psi_{\theta},\partial_{i}\partial_{j}\psi_{\theta}\rangle+\langle\partial_{i}\psi_{\theta},\psi_{\theta}\rangle\langle\psi_{\theta},\partial_{j}\psi_{\theta}\rangle\right\}$
@@
$\triangle F=\frac{-1}{2}$ $\left[F({\bm{\theta}}_{k}+s_{k}\vec{\Delta}^{1}_{k}+s_{k}\vec{\Delta}^{2}_{k},{\bm{\theta}}_{k})\right.$ (32)
@@
$-\left.F({\bm{\theta}}_{k}-s_{k}\vec{\Delta}^{1}_{k}+s_{k}\vec{\Delta}^{2}_{k},{\bm{\theta}}_{k})\right],$
which is composed of four terms respective to four quantum expectations we run on the quantum processor. Then, the Fubini-Study metric Eq. (30) is replaced by the QN-SPSA metric at $k$-th iteration
@@
where $\bar{H}^{k}({\bm{\theta}})\in\mathbb{R}^{p\times p}$ and $s_{k}$ is a small positive hyper-parameter being able to be tuned. The metric is, however, still too much stochastic, some helpful techniques are invoked to resolve that hitch [30, 31], say, counting information from previous updates to smooth the estimator
@@
eventually, because the estimator is still ill-conditioned and unstable. To satisfy the positive semi-definite warranting the local convex analysis and invertibility condition of $\tilde{H}^{k}$, the below replacement is needed accordingly,
@@
where the second term $\beta\mathbb{1}\in\mathbb{R}^{p\times p}$ corresponds to the invertibility condition. The effect from geometry metric is suppressed as the positive regulator gets a huge value $\beta\gg 0$, which indeed reduces to the standard gradient descent scenario and is more unstable. Therefore, that, in turn, causes a larger deviation in the average sample result when $\beta\rightarrow 0$. That makes a constant regulator $\beta$ is thus a trade-off between Quantum Natural information and numerical instability. Another regularization approach, called “half-inversion”, generalizes the power of the second-order derivative matrix to $n$ instead of the usual ones $-1$ corresponding to standard natural gradient and 00 standing for original gradient, given by [32].
@@
Quantum Natural-Simultaneous Perturbation Stochastic Approximation with Parameter-Shift Rule (QN-SPSA+PSR) is our extension for the QN-SPSA+SPSA method. Although QN-SPSA+SPSA has smaller complexity when compared with other methods, this stochastic method seems so randomized and made it harder to converge and less stable when it comes to a bigger cost function landscape. Here, We propose the QN-SPSA+PSR algorithm, which is inherent in approximating the Fubini-Study metric properties in a way of saving computational cost. However, instead of using the SPSA method to approximate the gradient of the cost function, we use the analytical method to calculate the gradient by PSR, which would compensate for the instability of the stochastic approximation in the Fubini-Study metric and drive itself in the gradient direction. This is expected for more stable and fast convergence compared to the former method QN-SPSA+SPSA.
V Results
V.1 Complexity estimation

### v2

@@ II.2 Transverse Ising model
The Transverse Ising model (TIM) was introduced first in 1963 by de Gennes when he built a model for describing the low-frequency collective modes of protons in the ferroelectric phase of ferroelectric crystals $\text{KH}_{2}\text{PO}_{4}$ [37]. From then until now, the Transverse Ising Model has been used to represent enormous simple-to-complex systems such as brain science and information science and technology.
@@
One of the key differences between the classical and quantum versions of the Ising model is that the expectation value of two terms in Eq. (4) must be measured separately due to the non-commutative feature of these elements. However, the commutation property allows us to perform only two observations to extract the Hamiltonian’s expectation. The first term $\sum_{n=1}^{N-1}\sigma^{z}_{n-1}\sigma^{z}_{n}$ can be cumulatively achieved by measuring directly on the computational basis as follows:
@@ II.3 Ansatz Construction
II.3.1 Symmetry of the Transverse Ising model
@@
  • Real representation. Using the eigenstates of the $\sigma_{Z}$ (Pauli-z) operator as the elementary binary computational basis, in terms of which, $\sigma^{z}$, $\sigma^{x}$ are real matrices, due to that, we can represent the TIM Hamiltonian in real form, the eigenstates of the TIM Hamiltonian can thus be chosen to be real for conventional purposes. Namely, considering $|\Psi\rangle=\sum_{n}C_{n}|n\rangle$ is an eigenstate of $H_{\text{TIM}}$ expanding in the computational basis $|n\rangle$. The real form of $H_{\text{TIM}}$ means the Hermitian real element matrix that induces coefficients satisfying $C^{*}_{n}C_{m}=C_{n}C^{*}_{m}$ $\forall m,n\in[0,2^{N}-1]$, then, in generality
@@
    The interaction of two local spins formulated by this term causes a sort of entanglement structure between every two local spins of the ground state energy. Nevertheless, in the case of the order phase, when the external magnetic field is dominant, this interaction can be broken down, each spin will be freely interacting and align in the same direction as the magnetic field.
  • Total spin-flip symmetry . The most featured symmetry of the Ising model can be referred to as the classical counterpart, time-reversal symmetry (generally, $\mathcal{Z}_{2}$ symmetry). Under the total spin-flip transformation in the z direction $\left(\sigma^{x}\right)^{\otimes N}$
@@
II.3.2 Ansatz selection
For the chosen ansatz, the conventional real coefficients of eigenstates tell us that it is enough to span in real quantum parameter space for finding the ground state energy, and the linear entanglement mapping comes from the information of the local interaction term in the Hamiltonian. Moreover, to accommodate the device constraints, in a speculative way, the available common-use RealAmplitudes ansatz is our good candidate for the implementation, which uses only the quantum gate $R_{y}\left(\bf\theta\right)$ for the parametrized quantum circuit and the entanglement scheme as in Fig. 2a.
[…]
The additional reason for choosing such a Hardware-efficient ansatz architecture [22] comes from the efficient implementation of quantum optimization to attain our full quantum algorithm purpose, which will be addressed in the next subsection III.2.
To efficiently travel around the full quantum Hilbert space, we need to tune $22^{N+1}-$ 2 degrees of freedom, $+1$ inside of the exponent indicates the complex number condition and $-2$ turns up from normalization and global phase conditions. After bringing TIM’s properties, we investigated from section II.3.1 to dissect. Real representation Eq. (7) helps us eliminate $+1$ within the exponent, and Total spin-flip symmetry Eq. (11) reduces the number of parameters we have to search the ground state by half. Finally,
@@ II.3.2 Ansatz selection
The number of layers $L$, which changes directly with the ansatz number of parameters $p$ according to $p=N(L+1)$. Following Eq. (12), we roughly evaluate a necessary number of layer
@@
notes that this is not a rigorous required condition because the ansatz’s entanglement mapping scheme used also contributes, which is able to change the bound as well, it is still a good estimation for surface analysis, nonetheless. Yet, we can realize that, not only TIM, but also Eq. (12) holds for all symmetrically real Hamiltonians being symmetric under $\left(\sigma^{x}\right)^{\otimes N}$ the operator. Therefore, there are some hidden TIM properties that we are unable to analytically incorporate to give an exact boundary value of $L$. Through conducting the experimental survey though, we are able to choose a reliable value, as we will see in chapter IV.2, it works well even for choosing $L=2$ for all qubit scaling.
III Optimization
The next subroutine in the VQE work map, where the optimization of variational parameters plays a key role in hybrid quantum-classical algorithms, and requires efficient methods to achieve fast and accurate convergence. The choice of optimization methods directly impacts the performance of the whole algorithm, influencing their ability to reach optimal solutions with minimal computational cost. Our goal is to investigate different optimization techniques, analyzing their performance to enhance computational efficiency and stability.
Effective optimization methods are essential not only for VQE but also for other variational approaches for quantum chemistry, material simulations and combinatorial optimization tasks. For broader applications, in Quantum Machine Learning (QML) that plays a critical role in training quantum neural networks, kernel-based models, and generative learning algorithms.
In this work, we focus on the Transverse Ising Model as a test bench, where we minimize the cost function defined as the expectation value of the Hamiltonian, $\langle\hat{H}_{\text{TIM}}\rangle$, using common methods of optimization categorized into two class operations.
III.1 Classical Operation
@@
Constrained Optimization BY Linear Approximation (COBYLA). One of the most powerful derivative-free methods, which has been favored by many users in lately decades, COBYLA makes use of linear interpolation of the objective function at each iteration by a unique linear polynomial function at the vertices for finding an optimal vector parameter within the trust region, then feeding the optimal point evaluated to the objective function to get the value improving the next iteration of approximation [38, 39, 40, 41]. In VQE, COBYLA’s linear interpolation at each step is a classical subroutine running on the CPU using only one objective value from the QPU evaluation running. And so, because of ignored derivative information, it would beneficially avoid several problems in analysis optimizations, especially Barren plateaus landscape, in return for less accuracy for more parameters ($p>9$) [41]. For a bird’s-eye view, new updates to COBYLA versions like UOBYQA, NEWUOA, and BOBYQA count the objective function’s curvature information to increase convergence [42, 43, 44].
Finite Difference (FD). Let’s dawn on one of the primary first-order derivative-based approaches, where the parameter ${\bm{\theta}}\in\mathbb{R}^{p}$ is updated at the $k$-th iteration by
@@
where $\vec{i}\in\mathbb{R}^{p}$ is $i$-th the unit vector and $\epsilon$ is an infinitesimal change, with the error proportional to $\mathcal{O}(\epsilon^{2})$. The smaller $\epsilon$ we use, the more exact the result we get. However, due to the restricted accuracy of a classical computer, we are unable to achieve a value of $\epsilon$ that is too small, which even becomes greater when implemented on a quantum device, where at least the sample error enters the picture.
Simultaneous Perturbation Stochastic Approximation (SPSA). To surmount some of the obstacles that arise from gradient descent operation on near-term devices, the idea of a perturbing stochastic approximation method is a good choice. SPSA generates an unbiased estimator $\tilde{f}({\bm{\theta}}_{k})$ of the gradient by simultaneously randomly perturbating the gradient direction of parameters [45], we replace the ordinary gradient vector $\triangledown f({\bm{\theta}}_{k})$ by
@@
where $\vec{\Delta}_{k}$ is a random perturbed vector sampled from the zero-mean distribution, usually, the Bernoulli distribution is used. We can see that all parameters are simultaneously shifted by a random amount (±$\pm s_{k}$), our computation, therefore, only requires two objective function evaluations per iteration. Whereas standard gradient computation time is scaled along with the number of parameters, the SPSA optimizer is, however, independent, which saves a lot of time as we work on a higher parameter regime. Besides that, noise from quantum circuit executions computing the objective value can be regarded as a stochastic perturbation part absorbed into the SPSA procedure. These advantages promote SPSA and its other versions to be efficient techniques in the NISQ era.
III.2 Quantum Operation
As opposed to the classical ones, the quantum operations require further quantum information extracted from the structure of the problem, such as the trial wave function. This sort of property inevitably entails the details throughout the problem analysis, especially regarding Ansatz Construction.
Parameter-shift rules (PSR). Inspired by the classical shift rules in the exact derivative computation of some special functions, we are successfully able to figure out the analytical value of the objective function’s derivatives using quantum devices. For any kind of quantum gate that has the form
@@ III.2 Quantum Operation
generated by the Hermitian operator $\hat{G}$. Suppose that we have an expectation value of $\hat{H}$, which is our objective function $f(\theta)=\langle\psi(\theta)|\hat{H}|\psi(\theta)\rangle$ parametrized by $\theta$. The ansatz wave function $|\psi(\theta)\rangle=\hat{U}(\theta)|\psi\rangle_{I}$ is made up of $\hat{U}(\theta)=\hat{A}\hat{\mathcal{G}}(\theta)\hat{B}$, with $\hat{A}$, $\hat{B}$ are other arbitrary operators. As a result, the partial derivative of the $f(\theta)$ with respect to $\theta$ is thus obtained via the parameter-shift rules [46]
@@
The partial derivative $\partial_{\theta}=\frac{\partial}{\partial\theta}$ implies we can generalize to a set of multiple parameters $\{\theta_{i}\}$. The evaluation of $f(\theta\pm\frac{\pi}{4s})$ can be easily run on quantum computers, and then the exact amplitude of the vector gradient is derived. Within the aim of this paper, the given parametric ansatz is made out of Pauli rotations, and $s=\frac{1}{2}$ is selected accordingly. The state-of-the-art formalism for more generic PQCs is also able to be derived [47, 48].
Quantum Natural Gradient Descent (QNG). How to improve the convergence? The global fixed learning rate $\eta_{k}$ is apparently not a good choice, as one does not be good sensitive to the model information with respect to parameter changes. Initially, many attempts try to tune the learning rate $\eta_{k}$, such as learning rate schedulers that vary $\eta_{k}$ after an iteration, latterly, adaptive methods that count previous iteration values or capture the curvature of the objective function using the diagonal approximation of the Hessian (HES), where each $\eta_{k}$ is the inverse Hessian’s diagonal element instead of being equally fixed. Those methods suggest that we can embed further information, rather than only the gradient vector, to make the optimal step size for our variational quantum algorithms. Similarly to the classical counterpart using the Fisher Information Matrix (FIM), Quantum Natural Gradient Descent invokes the quantum geometry of the wave function, which can help us get the optimal spot faster. Namely, we transform to the Riemann parameter space using the metric $g\in\mathbb{R}^{p\times p}$ of projected Hilbert space $\mathcal{PH}$ instead of the current Euclidean space $g=\mathbb{1}_{p\times p}$. The update Eq. (14) turns to
@@
We expand up $|\psi_{\theta+d\theta}\rangle=|\psi_{\theta}\rangle+|\partial_{i}\psi_{\theta}\rangle d\theta_{i}+\frac{1}{2}|\partial_{i}\partial_{j}\psi_{\theta}\rangle d\theta_{i}d\theta_{j}$ upto second order in $d\theta$, where $\partial_{i}=\frac{\partial}{\partial\theta_{i}}$. The Quantum Geometry tensor (QGT) $G_{ij}\in\mathbb{R}^{p\times p}$ has two parts: the anti-symmetric imaginary part $\sigma_{ij}=-\sigma_{ji}$ related to the gauge field of $U(1)$ eventually vanishes, the only contribution comes from the real symmetric part $g_{ij}\equiv g_{ij}({\bm{\theta}})$ reflects the quantum distance in $\mathcal{PH}$ space, where its formula is
@@
Quantum Natural Gradient - Block Diagonal Approximation (QN-BDA). Moreover, apart from Diagonal Approximation, where we just count $p$ elements in the diagonal line, the QN-BDA employing the Parametric Family circuit is conveniently deployed entirely in the computation of the approximated metric tensor on the quantum computer [50]. The parametric unitary operator $\hat{U}({\bm{\theta}})$ acts on the initial state, $|\psi\rangle_{I}$ entailing $L$ layers
@@
  • $S_{l}$ are the static parts, which usually denote the entanglement layer.
@@
  • The single qubit gate $R_{i}(\theta^{l}_{i})=\exp{\left[i\theta_{i}^{l}K_{i}\right]}$ ) is constructed from the Hermitian generator $K_{i}$ with the parameter $\theta_{i}^{l}\in\theta^{l}=\left\{\theta^{l}_{1},\ldots,\theta^{l}_{N}\right\}$.
Such a type of parametrized circuit, whose nice properties we can prospect to yield a block diagonal form of QGT running completely on the quantum processor, each block corresponding to each layered vector parameter $\theta^{l}\in{\bm{\theta}}=\theta^{1}\oplus\ldots\oplus\theta^{L}$. We denote
@@
$|\partial_{i}\psi_{\theta}\rangle$ $=\partial_{i}\hat{U}({\bm{\theta}})|\psi\rangle_{I}=\hat{U}^{L}_{l+1}S_{l}\partial_{i}P_{l}(\theta^{l})\hat{U}^{l-1}_{1}|\psi\rangle_{I}$ (25)
$=\hat{U}^{L}_{l}\left(iK_{i}\right)|\psi^{(l-1)}_{\theta}\rangle.$
Note that $\left[K_{i},K_{j}\right]=0$ in each layer, the unitarity of $\hat{U}^{m}_{n}$ and the overlap of partial derivative states tell us the block matrix element
@@
of the block-diagonal form of the Fubini-Study metric
@@
Quantum Natural-Simultaneous Perturbation Stochastic Approximation (QN-SPSA). Inheriting the SPSA idea to calculate the gradient, it is capable to put up generating the Hessian matrix with fewer evaluations, the so-called 2-SPSA. Normally, the HES or FIM matrix consumes $\mathcal{O}(p^{2})$ quantum expectation. QN-BDA enables us to turn to relying on the number of layers $\mathcal{O}(L)$, saving us lots of computational resources. Even further, QN-SPSA just only executes four quantum runs to obtain the full second-order derivative matrix. Let’s consider the Hessian of the Fubini-Study metric,
@@
which is just the Hessian form of Eq. (22) so that we can deploy the 2-SPSA method to generate the QN-SPSA matrix [51, 52]. We can see the equivalence,
@@
$\triangle F=\frac{-1}{2}$ $\left[F({\bm{\theta}}_{k}+s_{k}\vec{\Delta}^{1}_{k}+s_{k}\vec{\Delta}^{2}_{k},{\bm{\theta}}_{k})\right.$
@@
$-\left.F({\bm{\theta}}_{k}-s_{k}\vec{\Delta}^{1}_{k}+s_{k}\vec{\Delta}^{2}_{k},{\bm{\theta}}_{k})\right],$ (29)
which is composed of four terms, corresponding to four quantum expectations we run on the quantum processor. Then, the Fubini-Study metric Eq. (27) is replaced by the QN-SPSA metric at $k$-th iteration
@@
where $\bar{H}^{k}({\bm{\theta}})\in\mathbb{R}^{p\times p}$ and $s_{k}$ is a small positive hyperparameter able to be tuned. The metric is, however, still too highly stochastic, so some helpful techniques are invoked to resolve that hitch, say, counting information from previous updates to smooth the estimator
@@
eventually, because the estimator is still ill-conditioned and unstable. To satisfy the positive semi-definite warranting the local convex analysis and invertibility condition of $\tilde{H}^{k}$, the following replacement is needed accordingly,
@@
where the second term $\beta\mathbb{1}\in\mathbb{R}^{p\times p}$ corresponds to the invertibility condition. The effect from the geometry metric is suppressed as the positive regulator gets a huge value $\beta\gg 0$, which indeed reduces to the standard gradient descent scenario and is more unstable. Therefore, that, in turn, causes a larger deviation in the average sample result when $\beta\rightarrow 0$. That makes a constant regulator $\beta$, which is thus a trade-off between Quantum Natural information and numerical instability. Another regularization approach, called “half-inversion”, generalizes the power of the second-order derivative matrix to $n$ instead of the usual ones $-1$ corresponding to standard natural gradient and 00 standing for original gradient, given by [53].
@@
Quantum Natural-Simultaneous Perturbation Stochastic Approximation with Parameter-Shift Rule (QN-SPSA+PSR) is our extension for the QN-SPSA+SPSA method. Although QN-SPSA+SPSA has smaller complexity when compared with other methods, this stochastic method seems so randomized and makes it harder to converge and less stable when it comes to a bigger cost function landscape. Here, we propose the QN-SPSA+PSR algorithm, which is inherent in approximating the Fubini-Study metric properties in a way of saving computational cost. However, instead of using the SPSA method to approximate the gradient of the cost function, we use the analytical method to calculate the gradient by PSR, which would compensate for the instability of the stochastic approximation in the Fubini-Study metric and drive itself in the gradient direction. This is expected for more stable and faster convergence compared to the former QN-SPSA+SPSA method.
IV Results and Discussion
IV.1 Complexity estimation